OpenStudy (emilyjones284):

help please i give medals(:

OpenStudy (emilyjones284):

\[\sqrt{15p}\over \sqrt{5}\]

OpenStudy (emilyjones284):

simplify.

OpenStudy (anonymous):

Multiply top and bottom by the square root of 5: \[\frac{ \sqrt{15p} }{ \sqrt{5} } = \sqrt{ \frac{ 15p }{ 5 }} = \sqrt{3p}\]

OpenStudy (anonymous):

Sorry, it messed up my answer and I'm not sure why. Ignore the text above the equation. You need to combine it into one radical, as I did in the equation.

OpenStudy (emilyjones284):

okay but the answer would be square root of 3p?

OpenStudy (anonymous):

yep